Puget World, Inc., manufactures two models of television sets, the N 800 XL model and the N 500 model. Data regarding the two products follow:

Direct Labor-
Hours per Unit
Annual
Production
Total Direct
Labor-Hours
Model N 800 XL 3.0 4,000 units 12,000
Model N 500 1.0 13,000 units 13,000
25,000

Additional information about the company follows:

a. Model N 800 XL requires $70 in direct materials per unit, and Model N 500 requires $30.

b. The direct labor wage rate is $17 per hour.

c. The company has always used direct labor-hours as the base for applying manufacturing overhead cost to products.

d. Model N 800 XL is more complex to manufacture than Model N 500 and requires the use of special equipment. Consequently, the company is considering the use of activity-based costing to assign manufacturing overhead cost to products. Three activity cost pools have been identified as follows:

Activity Cost Pool Activity
Measure
Estimated
Overhead Cost
Machine setups Number of setups $ 520,000
Special processing Machine-hours 235,000
General factory Direct labor-hours 1,445,000
$ 2,200,000

Expected Activity

Activity Measure Model N 800 XL Model N 500 Total
Number of setups 150 250 400
Machine-hours 20,000 0 20,000
Direct labor-hours 12,000 13,000 25,000

Problem 4-17 Part 1

Required:

1. Assume that the company continues to use direct labor-hours as the base for applying overhead cost to products.

a. Compute the predetermined overhead rate.

b. Compute the unit product cost of each model.

Problem 4-17 Part 2

2. Assume that the company decides to use activity-based costing to assign manufacturing overhead cost to products.

a(1). Compute the activity rate for each activity cost pool.

a(2). Determine the amount of overhead cost that would be assigned to each model using the activity-based costing system.

b. Compute the unit product cost of each model.
Answer #1

Ansuky: @ compert the predetermined overhead rate : pocketermined roete per DLH = 2,300,000 25,000 = $88 por DLH ...predeterm= $30+ (1x17) + (1x88) = $38 + r7 +88 = $135 os model N 500 total cost = $135| 2-a) cal cellation de predetermined al locatio2 am model N800; Total cost of machine setups - Activity Driver x activity rate = 150 X 1300 = 195,000 Total cost do siecialmocede nu 500; Total cost, machine selups & activity Drivery attivity vete innvolürty karte = 250 X 1300 $325,000 processingmodel N soor Toted cost = Direet Merterials + Direct labort overladt S. = 30+ (1x 17) +83 - 130